OpenStudy (anonymous):

classify all critical points k(x, y) = x^(2) − xy + 4y^(2)

OpenStudy (anonymous):

gradient of k(x,y) is {2 x - y, -x + 8 y} See when it is equal to {0,0} and you get a critical point

OpenStudy (anonymous):

2 x - y=0 -x + 8 y=0 will get us x=y=0 So we have to study the critical point (0,0)

OpenStudy (anonymous):

that makes sense! thank you :)

OpenStudy (anonymous):

You have now to study the Hessian Determinant at (0,0)

OpenStudy (anonymous):

The hessian determinant is always 15 for this function so we need to study the second derivative

OpenStudy (anonymous):

The second derivative of k \(x,y0 with respect to x is always 2, so the point (0,0) is a minimum

OpenStudy (anonymous):

Here is a plot of your function